### v1

@@ 1 Introduction
Now we know explicit families of $0/1$-polytopes that do not admit any linear or positive semidefinite extension of polynomial size [8], [9] and [10]. Surprisingly, for several decades after the work of Yannakakis [7], it was an open question whether every $0/1$-polytope in $\mathbb{R}^{n}$ has linear extension complexity that is polynomial in $n$. For the first time this question was answered in negative by Rothvoss [11]. The arguments in the work of Rothvoss [11] are based on counting and can be starightforwardly extended to stronger collections of extensions, as long as these collections admit a normalization of factorizations. By providing a normalization of factorizations for positive semidefinite cones in [12], a negative answer was provided also for a stronger collection of extensions, so-called positive semidefinite extensions. Later another possible normalization for positive semidefinite cones was given in [6]. Our work is motivated by providing normalization results for more general convex cones than the nonnegative orthant and positive semidefinite cones and improving our understanding of normalizations for convex cones.
@@ 3 Normalization of Cones
    $\langle{-\nabla F(x)},{h}\rangle^{2}\leq\vartheta\langle{\nabla^{2}F(x)h},{h}\rangle,$ for all $x\in\operatorname{int}(\mathcal{C}),$ $\forall h\in\mathbb{R}^{n}$;
@@
    $\langle{\nabla^{2}F(x)h},{h}\rangle^{1/2}\leq\langle{-\nabla F(x)},{h}\rangle,$ for all $x\in\operatorname{int}(\mathcal{C}),$ $\forall h\in\mathcal{C}$;
@@
The fact that every regular cone in $\mathbb{R}^{n}$ admits a $\vartheta$-LHSCB $F$, where $1\leq\vartheta\leq n$, follows from Theorem 2.5.1 in [15] and [16]. The properties (a) and (b) are essentially parts of a definition of $\vartheta$-LHSCB. For part (c), see [15, Proposition 2.3.4, eqn. (2.3.12)]. For parts (e) see [17, eqn. (2.11)]. For part (d) see [18, Proposition 5.1]. For parts (f) and (g) see [19, Theorem 1.1 and eqns. (1.17 and (1.18)], as well as [20, Section 4]. ∎
@@ 3.1 Symmetric Cones
    $\nabla^{2}F(x),\left[\nabla^{2}F(x)\right]^{1/2}\in\operatorname{Aut}(\mathcal{C}),$ $\forall x\in\operatorname{int}(\mathcal{C})$;
@@ 3.2 Homogeneous Cones
For parts a, b and c see [29, Section 8 and Theorem 6.1]. For part d see [21, Theorem 5.3]. ∎
@@
The dual cone $\mathcal{C}^{\star}$ is not facially exposed and therefore is not homogeneous [30]. So the results in this section are not applicable to the cone $\mathcal{C}$. Note however that $\mathcal{C}$ is a hyperbolic cone (intersection of two well-known hyperbolic cones: a closed half-space and a three-dimensional second-order cone) and its dual $\mathcal{C}^{*}$ is not [31, Example 4].

### v2

@@ 1 Introduction
Now we know explicit families of $0/1$-polytopes that do not admit any linear or positive semidefinite extension of polynomial size [8], [9] and [10]. Surprisingly, for several decades after the work of Yannakakis [7], it was an open question whether every $0/1$-polytope in $\mathbb{R}^{n}$ has linear extension complexity that is polynomial in $n$. For the first time this question was answered in negative by Rothvoss [11]. The arguments in the work of Rothvoss [11] are based on counting and can be starightforwardly extended to stronger collections of extensions, as long as these collections admit a normalization of factorizations.
In their work on positive semidefinite cones, Briët et al. [12] provided a negative answer for a broader class of extensions known as positive semidefinite extensions by introducing a normalization approach. Subsequent research by Fawzi et al. [6] expanded on this by presenting another normalization method using John-Löwner ellipsoids. Averkov et al. [13] similarly employed these ellipsoids to analyze the semidefinite extension complexity of polytopes.
The utility of John-Löwner ellipsoids for normalization has proven versatile, finding applications beyond the analysis of positive semidefinite cone extension complexity, such as in Banach space theory, as mentioned in the work [14, Lemma 4.2]. Building on these foundations, our research aims to extend normalization results to more general convex cones, seeking to broaden our understanding of cone normalizations.
Interestingly, while our arguments in Section 3 may initially appear very distinct, they nonetheless share underlying connections to the problem of constructing structured minimum volume ellipsoids that contain specific sets.
@@ 3 Normalization of Cones
    $\langle{-\nabla F(x)},{h}\rangle^{2}\leq\vartheta\langle{\nabla^{2}F(x)h},{h}\rangle,$ for all $x\in\operatorname{int}(\mathcal{C})$ and for all $h\in\mathbb{R}^{n}$;
@@
    $\langle{\nabla^{2}F(x)h},{h}\rangle^{1/2}\leq\langle{-\nabla F(x)},{h}\rangle,$ for all $x\in\operatorname{int}(\mathcal{C})$ and for all $h\in\mathcal{C}$;
@@
The fact that every regular cone in $\mathbb{R}^{n}$ admits a $\vartheta$-LHSCB $F$, where $1\leq\vartheta\leq n$, follows from Theorem 2.5.1 in [17] and [18]. The properties (a) and (b) are essentially parts of a definition of $\vartheta$-LHSCB. For part (c), see [17, Proposition 2.3.4, eqn. (2.3.12)]. For parts (e) see [19, eqn. (2.11)]. For part (d) see [20, Proposition 5.1]. For parts (f) and (g) see [21, Theorem 1.1 and eqns. (1.17 and (1.18)], as well as [22, Section 4]. ∎
@@ 3.1 Symmetric Cones
    $\nabla^{2}F(x),\left[\nabla^{2}F(x)\right]^{1/2}\in\operatorname{Aut}(\mathcal{C})$ for all $x\in\operatorname{int}(\mathcal{C})$;
@@ 3.2 Homogeneous Cones
For parts a, b and c see [31, Section 8 and Theorem 6.1]. For part d see [23, Theorem 5.3]. ∎
@@
The dual cone $\mathcal{C}^{\star}$ is not facially exposed and therefore is not homogeneous [32]. So the results in this section are not applicable to the cone $\mathcal{C}$. Note however that $\mathcal{C}$ is a hyperbolic cone (intersection of two well-known hyperbolic cones: a closed half-space and a three-dimensional second-order cone) and its dual $\mathcal{C}^{*}$ is not [33, Example 4].
@@
Building upon the literature, we reference Remark 2 by Averkov et al. [13], which highlighted the critical role of automorphisms of the positive semidefinite cone in their proof. Our results in the preceding and current sections further demonstrate that automorphisms continue to play a significant role in more general convex cone settings.